# Amortized Variational Transdimensional Inference

### v1

@@ Abstract
The expressiveness of flow-based models combined with stochastic variational inference (SVI) has, in recent years, expanded the application of optimization-based Bayesian inference to include problems with complex data relationships. However, until now, SVI using flow-based models has been limited to problems of fixed dimension. We introduce CoSMIC  normalizing flows (COntextually-Specified Masking for Identity-mapped Components), an extension to neural autoregressive conditional normalizing flow architectures that enables using a single amortized variational density for inference over a transdimensional target distribution. We propose a combined stochastic variational transdimensional inference (VTI) approach to training CoSMIC  flows using techniques from Bayesian optimization and Monte Carlo gradient estimation. Numerical experiments demonstrate the performance of VTI on challenging problems that scale to high-cardinality model spaces.
@@ 1 Introduction
Amortized variational inference (Dayan et al., , 1995) is a universal topic in statistics and machine learning that has seen a surge in interest since the introduction of normalizing flows (Rezende and Mohamed, , 2015). An amortized density function can be used for a variety of downstream tasks, such as importance sampling (Rezende and Mohamed, , 2015), simulation-based inference (Papamakarios et al., , 2019; Zammit-Mangion et al., , 2024), adaptive MCMC Gabrié et al., (2022), and generative modeling (Kingma and Welling, , 2013). While many existing approaches only consider continuous supports, there is a growing interest in applications where the support is either discrete or discretely indexed (Diluvi et al., , 2024). One such application concerns a target transdimensional probability distribution $\pi$ with support
$\pazocal{X}$ $=\bigcup_{m\in\pazocal{M}}(\{m\}\times\Theta_{m}),$ (1)
where $\pazocal{M}$ is a finite discrete index set, $\Theta_{m}\subseteq\mathbb{R}^{d_{m}}$, and where the dimension $d_{m}$ of $\Theta_{m}$ may vary with $m$. Hence $\pazocal{X}$ is a transdimensional space (Green, , 1995; Sisson, , 2005). Such spaces arise naturally in Bayesian model inference, where the sets $\Theta_{m}$ correspond to model parameters, and $m\in\pazocal{M}$ is a model index. Model indices can parameterize various practical inference problems, such as variable selection, mixtures-of-regressions, or learning directed acyclic graphs (DAGs) from data.
This article is concerned with estimating the target distribution $\pi$ with associated density function $\pi(m,\bm{\theta}_{m})$, $\bm{\theta}_{m}\in\Theta_{m}$, which varies in dimension depending on $m$. (Note that for simplicity we refer to $\pi(m,\bm{\theta}_{m})$ and related functions as density functions, even though they are not continuous.) Typically, this density is only available in a conditional unnormalized form, $\eta(\bm{\theta}_{m}\,|\,m)=\pazocal{Z}_{m}\pi(\bm{\theta}_{m}\,|\,m)$, where $\pazocal{Z}_{m}=\int_{\Theta_{m}}\eta(\bm{\theta}_{m}\,|\,m)d\bm{\theta}_{m}$. The factorization $\eta(m,\bm{\theta}_{m})=\eta(\bm{\theta}_{m}\,|\,m)\pi(m)$ implies there is a discrete target probability mass function over models, $\pi(m)=\pazocal{Z}_{m}\pazocal{Z}^{-1}$, where $\pazocal{Z}=\sum_{m\in\pazocal{M}}\pazocal{Z}_{m}$. Estimation of $\eta(m,\bm{\theta}_{m})$ then becomes estimation of both $\eta(\bm{\theta}_{m}\mid m)$ and $\pi(m)$.
@@
In the presence of a likelihood function $p(\pazocal{D}\,|\,m,\bm{\theta}_{m})$ for data $\pazocal{D}$, and priors $p(\bm{\theta}_{m}\mid m)$ and $p(m)$, the target distribution is defined by the $\pazocal{D}$-conditional transdimensional posterior $\pi(m,\bm{\theta}_{m}\,|\,\pazocal{D})\propto p(\pazocal{D}\,|\,m,\bm{\theta}_{m})p(\bm{\theta}_{m}|m)p(m)$. In the context of variational Bayesian inference (Jordan et al., , 1999; Blei et al., , 2017), the process by which a variational approximation of the transdimensional posterior $\pi(m,\bm{\theta}_{m}\,|\,\pazocal{D})$ can be constructed has not been addressed in generality. Such a scheme would approximate some unnormalized target density $\eta(m,\bm{\theta}_{m}\,|\,\pazocal{D})=\pazocal{Z}\pi(m,\bm{\theta}_{m}\,|\,\pazocal{D})$ by optimizing over the parameters $\phi\in\mathbb{R}^{n_{\phi}},\psi\in\mathbb{R}^{n_{\psi}}$ of a tractable variational density family $q_{\psi,\phi}(m,\bm{\theta}_{m})=q_{\phi}(\bm{\theta}_{m}\mid m)q_{\psi}(m)$ via an objective function
$\psi^{*},\phi^{*}:=\operatorname*{arg\,min}_{\psi,\phi}\pazocal{L}(\psi,\phi),\qquad\pazocal{L}(\psi,\phi)=D_{\mathrm{KL}}(q_{\psi,\phi}||\pi)$ (2)
@@
where the loss $\pazocal{L}$ is the Kullback-Leibler (KL) divergence. There are two immediate impediments to constructing such a variational approximation. The first is defining and optimizing $q_{\phi}$ as $\bm{\theta}_{m}$ may vary in dimension conditional on $m$. The second is the inference of $q_{\psi}$ for discrete latent variables $m$ during the optimization of $q_{\phi}$, which amounts to optimization over non-stationary rewards as the convergence of $\phi\rightarrow\phi^{*}$ and $\psi\rightarrow\psi^{*}$ are interdependent.
@@
The MADE autoregressive encoder (Germain et al., , 2015) enabled autoregressive neural flow architectures, which can be coupled with any $\tau$ such as affine (Papamakarios et al., , 2017) and spline (Durkan et al., , 2019) based transformations. The computational complexity of the autoregressive flow is direction-dependent, being sequentially $O(d)$ in the inverse direction when compared with the forward direction. The inverse autoregressive flow (IAF) (Kingma et al., , 2016) avoids the sequential computational complexity in the variational setting by setting $\bm{\theta}=T_{\phi}(\bm{z})$, yielding the variational density $q_{\phi}(\bm{\theta}_{m})=\nu_{d}(T_{\phi}^{-1}(\bm{\theta}))|\det\nabla T_{\phi}^{-1}(\bm{\theta})|=\nu_{d}(\bm{z})|\det\nabla T_{\phi}(\bm{z})|^{-1}$.
Contributions: In this article, we introduce CoSMIC  (COntextually-Specified Masking for Identity-mapped Component) flows, a widely applicable and simple modification to conditional neural flow architectures (Section 2). CoSMIC  flows fundamentally expand the use cases for normalizing flows to encompass amortized variational inference applications so that a single amortized variational density can be used for inference over a transdimensional target distribution. In Section 3, we demonstrate the efficacy of CoSMIC  transformations within a novel variational transdimensional inference (VTI) framework with two implementations. The first builds upon principles of Bayesian optimization (Srinivas et al., , 2012), and the second uses Monte Carlo gradient estimation (Mohamed et al., , 2020). We also provide a theoretical analysis of VTI approximation error bounds under a Gaussian process surrogate, and convergence guarantees for the marginal model distribution under convergent optimization steps. Finally, we demonstrate the applicability of VTI to problems with model spaces that cannot be easily enumerated within the memory limitations of current computing architectures. In particular, Section 5 explores problems in Bayesian robust variable selection (O’Hara and Sillanpää, , 2009) and Bayesian model discovery of directed acyclic graphs (Heckerman et al., , 2006) (DAGs). Python/Pytorch CUDA code for all experiments will be made available at https://github.com/daviesl/avti.
@@ 2 Formulating a transdimensional variational density
The goal is to show that equation 5 factorizes into an active part and an i.i.d. auxiliary part, i.e.
@@
and to exploit this factorization in the construction of a transdimenisonal loss function. To achieve this factorization, we define the following notation. Let $A_{i}:\pazocal{M}\!\to\!\{0,1\}$ flag whether latent coordinate $i$ appears in model $m$, and let $B_{i}:\{0,1\}\!\to\!\{0,1\}^{|\bm{\rho}_{i}|},\;B_{i}(b)=(b,\ldots,b)$, broadcast this bit to the corresponding parameter block. Their composition $C_{i}:=B_{i}\circ A_{i}:\pazocal{M}\!\to\!\{0,1\}^{|\bm{\rho}_{i}|}$ therefore activates exactly the autoregressive parameters $\bm{\rho}_{i}$ needed by $\tau_{\bm{\rho}_{i}}(\bm{z}^{(i)})$ under model $m$. Concatenating the blocks gives the global context-to-mask map (see Figure 1(b) for a visualization):
@@
Similarly, $A$ and $B$ denote the respective coordinate-concatenated maps similar in form to equation 7. After a fixed left–right permutation aligning latents with $\bm{\theta}_{m}$, Proposition 2.2 proves this factorization is exact for any autoregressive network $\operatorname{NN}_{\phi}$ that parametrizes the transport $T_{\phi}$.
Recalling the univariate bijective maps of the inverse autoregressive flow as $\tau_{\bm{\rho}_{i}}:\mathbb{R}\mapsto\Theta_{i}$ for $i=1,\dots,{d_{\mathrm{max}}}$, we assume the existence of a static point $\bm{\rho}^{\mathrm{Id}}$ such that $\tau_{\bm{\rho}^{\mathrm{Id}}}(z)=z$ for all $z\in\mathbb{R}$, i.e., the transform becomes the identity map at $\bm{\rho}^{\mathrm{Id}}$. For example, a simple affine transformation (scale and location shift) is $\theta=\tau_{\bm{\rho}_{i}}(z)=\rho^{(0)}+\rho^{(1)}z$, where $\bm{\rho}_{i}=(\rho^{(0)},\rho^{(1)})$. In this case, the static point is $\bm{\rho}^{\mathrm{Id}}=(0,1)$ as then $\theta=z$. We can then construct simple mechanism for “choosing” between $\bm{\rho}_{i}$ and $\bm{\rho}^{\mathrm{Id}}$ for each individual transform $\tau$, $i=1,\dots,{d_{\mathrm{max}}}$, via the convex combination
@@
Each coordinate-wise transform then becomes $\bm{\theta}_{m}^{(i)}=\tau_{\bm{\rho}_{i}^{C}}(\bm{z}^{(i)})$, $i\in\{1,\dots,{d_{\mathrm{max}}}\}$. That is, the transformation parameters become a context-dependent composition of the elements of $\bm{\rho}_{i}$ and the static point $\bm{\rho}^{\mathrm{Id}}$ (Figure 1(c)). A composition of transforms parametrized according to equation 8 is a Context-Specified Masking for Identity-mapped Components (CoSMIC) normalizing flow.
@@
For a CoSMIC normalizing flow, $\bm{u}_{\setminus m}=\bm{z}_{\setminus m}$.
Proof. Let $I(m)=\{\,i\in\{1,\dots,{d_{\mathrm{max}}}\}:A_{i}(m)=1\,\}$, and $I^{c}$ be the complement. The result holds from equation 8 as, for all coordinates $i\in I^{c}(m)$, $\bm{u}_{\setminus m}^{(i)}=\tau_{\bm{\rho}_{i}^{C}}(\bm{z}_{\setminus m}^{(i)})=\tau_{\bm{\rho}^{\mathrm{Id}}}(\bm{z}_{\setminus m}^{(i)})=\bm{z}_{\setminus m}^{(i)}.\hskip 28.45274pt\square$
Proofs of the following Proposition 2.2 and Corollary 2.3 are in Appendix˜A.
@@
Fix $m\in\pazocal{M}$. Let $P_{m}$ be the permutation matrix that places the coordinates indexed by $I(m)$ (from Lemma˜2.1) before those in $I^{c}(m)$ while preserving the original order inside each group. Define the left-right-permuted flow $T_{\phi}^{(P_{m})}:=P_{m}^{-1}\!\circ T_{\phi}\circ P_{m}$ and the corresponding density $\tilde{q}_{\phi}^{(P_{m})}(\bm{\theta}_{m},\bm{u}_{\setminus m})=\nu_{{d_{\mathrm{max}}}}(\bm{z}))\,\bigl|\det\nabla T_{\phi}^{(P_{m})}(\bm{z}\mid m)\bigr|^{-1}$, $\bm{z}=T_{\phi}^{(P_{m}),-1}(\bm{\theta}_{m},\bm{u}_{\setminus m})$. Redefine $C:=C^{(P_{m})}=B\circ P_{m}\circ A$. Then (a) $\tilde{q}_{\phi}^{(P_{m})}(\bm{\theta}_{m},\bm{u}_{\setminus m})$ factorizes as per equation 6 with the substitution $\tilde{q}_{\phi}:=\tilde{q}_{\phi}^{(P_{m})}$, and (b) the marginal $q_{\phi}(\bm{\theta}_{m}\mid m)$ is consistent.
@@
and the loss function in equation 2 becomes
$\pazocal{L}(\psi,\phi)$ $=\mathbb{E}_{m\sim q_{\psi}}\left[\ell(m;\phi)-\log p(m)+\log q_{\psi}(m)\right],$ (10)
where
$\ell(m;\phi):=\mathbb{E}_{\bm{z}\sim\nu_{d_{\mathrm{max}}}}\left[\log h_{\phi}(\bm{z}\mid m)\right].$ (11)
The implementation of a CoSMIC inverse autoregressive flow step $T_{i}$ as part of a composition of transforms $T_{L}\circ\dots\circ T_{1}$ is visualized in Figure 1(a). Individual architectures for affine and rational quadratic spline transforms (Durkan et al., , 2019) and compositions are described in Section˜A.2.
[…]
Figure 1: Overview of (a) CoSMIC flow composition, (b) Context-to-mask map, (c) A single CoSMIC IAF step.

### v2

@@ Abstract
The expressiveness of flow-based models combined with stochastic variational inference (SVI) has expanded the application of optimization-based Bayesian inference to highly complex problems. However, despite the importance of multi-model Bayesian inference, defined over a transdimensional joint model and parameter space, flow-based SVI has been limited to problems defined over a fixed-dimensional parameter space. We introduce CoSMIC  normalizing flows (COntextually-Specified Masking for Identity-mapped Components), an extension to neural autoregressive conditional normalizing flow architectures that enables use of a single amortized variational density for inference over a transdimensional (multi-model) conditional target distribution. We propose a combined stochastic variational transdimensional inference (VTI) approach to training CoSMIC  flows using ideas from Bayesian optimization and Monte Carlo gradient estimation. Numerical experiments show the performance of VTI on challenging problems that scale to high-cardinality model spaces.
@@ 1 Introduction
Amortized variational inference [Dayan et al., 1995] has seen a surge in interest since the introduction of normalizing flows [Rezende and Mohamed, 2015]. Amortized densities can be used for a variety of downstream tasks, such as importance sampling [Rezende and Mohamed, 2015], simulation-based inference [Papamakarios et al., 2019, Zammit-Mangion et al., 2024], adaptive Markov chain Monte Carlo (MCMC) [Gabrié et al., 2022], and generative modeling [Kingma and Welling, 2013]. While many existing approaches only consider continuous supports, there is a growing interest in applications where the support is either discrete or discretely indexed [Diluvi et al., 2024]. One such application concerns a target transdimensional probability distribution $\pi$ with support $\pazocal{X}=\smash{\mathop{\textstyle\bigcup}\nolimits_{m\in\pazocal{M}}(\{m\}\times\Theta_{m})}$, where $\pazocal{M}$ is a finite discrete index set, $\Theta_{m}\subseteq\mathbb{R}^{d_{m}}$, and the dimension $d_{m}$ of $\Theta_{m}$ may vary with $m$. Hence $\pazocal{X}$ is a transdimensional space [Green, 1995, Sisson, 2005, Fan et al., 2026]. Such spaces arise in Bayesian model inference, where $\Theta_{m}$ correspond to model parameters, and $m\in\pazocal{M}$ is a model index. Discrete indices parameterize many practical inference problems, including variable selection [Fan et al., 2026], mixtures-of-regressions, learning directed acyclic graphs (DAGs) from data [Thompson et al., 2025], phylogenetic tree topology search Everitt et al. [2020], mixture-component inference Das and Bhattacharya [2014], geoscientific inversion Sambridge et al. [2013], and change-point models Green [1995]. This article is concerned with estimating the target distribution $\pi$ with associated density function $\pi(m,\bm{\theta}_{m})$, $\bm{\theta}_{m}\in\Theta_{m}$, whose dimension depends on $m$. For simplicity we refer to $\pi(m,\bm{\theta}_{m})$ and related functions as density functions, even though they are not continuous. Typically, this density is only available in a conditional unnormalized form, $\eta(\bm{\theta}_{m}\,|\,m)=\pazocal{Z}_{m}\pi(\bm{\theta}_{m}\,|\,m)$, where $\pazocal{Z}_{m}=\smash{\int_{\Theta_{m}}\eta(\bm{\theta}_{m}\,|\,m)d\bm{\theta}_{m}}$. The factorization $\eta(m,\bm{\theta}_{m})=\eta(\bm{\theta}_{m}\,|\,m)\pi(m)$ implies there is a discrete target probability mass function over models, $\pi(m)=\pazocal{Z}_{m}\pazocal{Z}^{-1}$, where $\pazocal{Z}=\smash{\sum_{m\in\pazocal{M}}\pazocal{Z}_{m}}$. Estimation of $\eta(m,\bm{\theta}_{m})$ then becomes estimation of both $\eta(\bm{\theta}_{m}\mid m)$ and $\pi(m)$.
@@
In the presence of a likelihood function $p(\pazocal{D}\,|\,m,\bm{\theta}_{m})$ for data $\pazocal{D}$, and priors $p(\bm{\theta}_{m}\mid m)$ and $p(m)$, the target distribution is defined by the $\pazocal{D}$-conditional transdimensional posterior $\pi(m,\bm{\theta}_{m}\,|\,\pazocal{D})\propto p(\pazocal{D}\,|\,m,\bm{\theta}_{m})p(\bm{\theta}_{m}|m)p(m)$. In the context of variational Bayesian inference (VI; see [Jordan et al., 1999, Blei et al., 2017]) approximation of the transdimensional posterior $\pi(m,\bm{\theta}_{m}\,|\,\pazocal{D})$ has not been addressed in generality. Such a scheme would approximate some unnormalized target density $\eta(m,\bm{\theta}_{m}\,|\,\pazocal{D})=\pazocal{Z}\pi(m,\bm{\theta}_{m}\,|\,\pazocal{D})$ by choosing parameters $\phi\in\mathbb{R}^{n_{\phi}},\psi\in\mathbb{R}^{n_{\psi}}$ of a tractable variational density family $q_{\psi,\phi}(m,\bm{\theta}_{m})=q_{\phi}(\bm{\theta}_{m}\mid m)q_{\psi}(m)$ to minimize
$\psi^{*},\phi^{*}:=\operatorname*{arg\,min}_{\psi,\phi}\pazocal{L}(\psi,\phi),\qquad\pazocal{L}(\psi,\phi)=D_{\mathrm{KL}}(q_{\psi,\phi}||\eta),$ (1)
@@
where $D_{\mathrm{KL}}$ is the Kullback-Leibler (KL) divergence. There are two impediments to constructing such a variational approximation: (i) defining and optimizing $q_{\phi}$ as $\bm{\theta}_{m}$ may vary in dimension conditional on $m$, and (ii) the inference of $q_{\psi}$ for discrete latent variables $m$ during the optimization of $q_{\phi}$, a non-stationary objective as $\phi\rightarrow\phi^{*}$ and $\psi\rightarrow\psi^{*}$ are interdependent.
@@
The MADE encoder [Germain et al., 2015] enables autoregressive neural flow architectures, which can be coupled with any $\tau$ such as affine [Papamakarios et al., 2017] and spline [Durkan et al., 2019] transformations. The cost of an autoregressive flow depends on the direction. In the forward (sampling) direction, it evaluates each dimension sequentially, for a time complexity of $\mathcal{O}(d)$. In the the inverse (likelihood) direction, computation can be parallel. The inverse autoregressive flow (IAF) [Kingma et al., 2016] reverses this dependence, setting $\bm{\theta}=T_{\phi}(\bm{z})$, yielding the variational density $\smash{q_{\phi}(\bm{\theta}_{m})=\nu_{d}(T_{\phi}^{-1}(\bm{\theta}))|\det\nabla T_{\phi}^{-1}(\bm{\theta})|=\nu_{d}(\bm{z})|\det\nabla T_{\phi}(\bm{z})|^{-1}}$.
Contributions: We introduce CoSMIC  (COntextually-Specified Masking for Identity-mapped Components) flows, a widely applicable and simple modification to conditional neural flow architectures (Section 2). CoSMIC  flows fundamentally expand the use cases for normalizing flows to encompass amortized variational inference applications, so that a single amortized variational density can be used for variational inference over a transdimensional (multi-model) target distribution. In effect, this extends the reparameterization trick exploited by IAF-based VI to the transdimensional setting. In Section 3, we demonstrate the efficacy of CoSMIC  transformations within a novel variational transdimensional inference (VTI) framework with two implementations. The first builds upon principles of Bayesian optimization [Srinivas et al., 2012], and the second uses Monte Carlo gradient estimation [Mohamed et al., 2020]. We also provide a theoretical analysis of VTI approximation error bounds under a Gaussian process surrogate, and convergence guarantees for the marginal model distribution under convergent optimization steps. Finally, we demonstrate the applicability of VTI to problems with model spaces that cannot be easily enumerated within the memory limitations of current computing architectures. In particular, Section 5 explores problems in Bayesian robust variable selection [O’Hara and Sillanpää, 2009] and Bayesian causal discovery [Heckerman et al., 2006].  [^1] PyTorch CUDA code for all experiments is available at https://github.com/daviesl/avti.
@@ 2 Formulating a transdimensional variational density
Our goal is to construct the IAF so that equation 4 factorizes into active and i.i.d. auxiliary parts, i.e.
@@
and to exploit this factorization in the construction of a transdimensional loss function. To achieve this factorization, we define the following notation. Let $A_{i}:\pazocal{M}\!\to\!\{0,1\}$ flag whether latent coordinate $i$ appears in model $m$, and let $B_{i}:\{0,1\}\!\to\!\{0,1\}^{|\bm{\rho}_{i}|},\;B_{i}(b)=(b,\ldots,b)$, broadcast this bit to the corresponding parameter block. Their composition $C_{i}:=B_{i}\circ A_{i}:\pazocal{M}\!\to\!\{0,1\}^{|\bm{\rho}_{i}|}$ therefore activates exactly the autoregressive parameters $\bm{\rho}_{i}$ needed by $\tau_{\bm{\rho}_{i}}(\bm{z}^{(i)})$ under model $m$. Concatenating the blocks gives the global context-to-mask map (see Figure 1(b) for a visualization):
@@
Similarly, $A$ and $B$ denote the respective coordinate-concatenated maps similar in form to equation 6. After a fixed left–align permutation aligning latents with $\bm{\theta}_{m}$, Proposition 2.2 proves this factorization is exact for any autoregressive network $\operatorname{NN}_{\phi}$ that parametrizes the transport $T_{\phi}$.
Recalling the univariate bijective maps of the inverse autoregressive flow as $\smash{\tau_{\bm{\rho}_{i}}:\mathbb{R}\mapsto\Theta_{i}}$ for $\smash{i=1,\dots,{d_{\mathrm{max}}}}$, we assume the existence of a static point $\smash{\bm{\rho}^{\mathrm{Id}}}$ such that $\smash{\tau_{\bm{\rho}^{\mathrm{Id}}}(z)=z}$ for all $z\in\mathbb{R}$, i.e., the transform becomes the identity map at $\bm{\rho}^{\mathrm{Id}}$. For example, a simple affine transformation (scale and location shift) is $\smash{\theta=\tau_{\bm{\rho}_{i}}(z)=\rho^{(0)}+\rho^{(1)}z}$, where $\smash{\bm{\rho}_{i}=(\rho^{(0)},\rho^{(1)})}$. In this case, the static point is $\smash{\bm{\rho}^{\mathrm{Id}}=(0,1)}$ as then $\theta=z$. We can then construct a simple mechanism for “choosing” between $\bm{\rho}_{i}$ and $\bm{\rho}^{\mathrm{Id}}$ for each individual transform $\tau$, $i=1,\dots,{d_{\mathrm{max}}}$, via the convex combination
@@
Each coordinate-wise transform then becomes $\smash{\bm{\theta}_{m}^{(i)}=\tau_{\bm{\rho}_{i}^{C}}(\bm{z}^{(i)})}$, $\smash{i\in\{1,\dots,{d_{\mathrm{max}}}\}}$. That is, the transformation parameters become a context-dependent composition of the elements of $\bm{\rho}_{i}$ and the static point $\bm{\rho}^{\mathrm{Id}}$ (Figure 1(c)). A composition of transforms parametrized according to equation 7 is a Contextually-Specified Masking for Identity-mapped Components (CoSMIC) normalizing flow.
@@
For a CoSMIC transform $(\bm{\theta}_{m},\bm{u}_{\setminus m})=T_{\phi}(\bm{z}_{m},\bm{z}_{\setminus m})$, $\bm{u}_{\setminus m}=\bm{z}_{\setminus m}$ $\forall m\in\pazocal{M}$.
@@
Fix $m\in\pazocal{M}$. Let $P_{m}$ be the permutation matrix that places the coordinates indexed by $I(m)$ (from the proof of Lemma˜2.1) before those in $I^{c}(m)$ while preserving the original order inside each group. Define the left-align-permuted flow $T_{\phi}^{\triangleleft}:=P_{m}^{-1}\!\circ T_{\phi}\circ P_{m}$ and the corresponding density $\smash{\tilde{q}_{\phi}^{\triangleleft}(\bm{\theta}_{m},\bm{u}_{\setminus m})=\nu_{{d_{\mathrm{max}}}}(\bm{z})\,|\det\nabla T_{\phi}^{\triangleleft}(\bm{z}\mid m)|^{-1}}$, $\smash{\bm{z}=T_{\phi}^{\triangleleft,-1}(\bm{\theta}_{m},\bm{u}_{\setminus m})}$. Redefine $C:=C^{\triangleleft}=B\circ P_{m}\circ A$. Then (a) $\tilde{q}_{\phi}^{\triangleleft}(\bm{\theta}_{m},\bm{u}_{\setminus m})$ factorizes as per equation 5 with the substitution $\tilde{q}_{\phi}:=\tilde{q}_{\phi}^{\triangleleft}$, and (b) the marginal $q_{\phi}(\bm{\theta}_{m}\mid m)$ is consistent.
@@
and, substituting $\ell(m;\phi):=\mathbb{E}_{\bm{z}\sim\nu_{d_{\mathrm{max}}}}\left[\log h_{\phi}(\bm{z}\mid m)\right]$, the loss in equation 1 becomes
$\pazocal{L}(\psi,\phi)$ $=\mathbb{E}_{m\sim q_{\psi}}\left[\ell(m;\phi)-\log p(m)+\log q_{\psi}(m)\right].$ (9)
Proposition 2.2 states that, conditional on model $m$, the CoSMIC IAF (Figure 1) achieves the factorization of the saturated-space variational approximation in equation 5. From Corollary 2.3, this means that when computing the loss function in equation 1, the ratio of the dimension-saturated variational density $\tilde{q}$ and conditional target $\tilde{\eta}$ (equation 8, LHS) which are both $d_{\mathrm{max}}$}-dimensional and which involve the auxiliary variables, collapses down to a direct comparison only on the $d_{m}$-dimensional model specific densities $q_{\phi}(\bm{\theta}_{m}\mid m)$ and $\eta(\bm{\theta}_{m}\mid m)$ (equation 8, RHS), and without the involvement of any auxiliary variables. That is, the CoSMIC flow enables the IAF to calculate on a fixed-dimensional space, while permitting the model-specific comparison within the loss function to operate on the natural $d_{m}$-dimensional space.
The implementation of a CoSMIC inverse autoregressive flow step $T_{i}$ as part of a composition of transforms $T_{L}\circ\dots\circ T_{1}$ is visualized in Figure 1(a). Individual architectures for affine and rational quadratic spline transforms [Durkan et al., 2019] and compositions are described in Appendix A.2.
Figure 1: (a) CoSMIC flow composition, (b) Context-to-mask map, (c) A single CoSMIC IAF step.